OpenStudy (anonymous):

x/x-1-3x/x^2+x-2=5/x+2 state the restrictions???

OpenStudy (anonymous):

could you add some brackets in order to make the question clearer?

OpenStudy (anonymous):

(x/x-1)-(3x/x^2+x-2)=5/x+2

OpenStudy (anonymous):

denominators cannot be zero so x^2+x-2=(x+2)(x-1) cannot =0 x-1 cannot =0 x+2 cannot =0 x can not be 1,-2

OpenStudy (anonymous):

thank you...so you factor out?

OpenStudy (anonymous):

factor the denominator in order to find out what values of x will make it zero